If 5sina=12cosa & a is an acute angle what is the value of cosec a???

Grade:10

1 Answers

Arun
6 years ago
Dear Shubham
 
As 
5 sinA = 12 cos A
Hence
sun A/ cos A = 12/5
tan A= 12/5
 
Hence sin A = 12/13
 
Now
cosec A = 1 /sin A = 13/12
 
Regards
Arun (askIITians forum expert)